Prove that an open interval in R is an open set and a closed intervals is a closed set
1
Expert's answer
2020-05-15T18:02:39-0400

Let's denote "A=(a,b)" and "M=[a,b]" are any arbitrary open and closed interval of "\mathbb{R}" respectively.

Claim: "A" is open set in "\mathbb{R}"

Proof:

We will show that every point of "A" is interior point of "A",hence "A" will be open in "\mathbb{R}" . Consider any arbitrary point "p \in A" and "r_p>0" , such that "B(p,r_p):=\{x \in \mathbb{R} |\: |x-p|<r_p\} \: (\clubs)" is the open "r_p" -ball of "p" ,Clearly if we choose suitable small enough "r_p" such "r_p" -ball always exist then "B(p,r_p) \subset A" ,Hence we are done.

Alternatively, observe that

"A=\cup_{p \in A} \: B(p,r_p) \: (\spades)"

Since, by the hypothesis,"B(p,r_p)" is open set , therefore union of open set is open.Hence "A" is open.


Now, consider the set "M", we claim that "M" is closed in "\mathbb{R}" .

Consider the compliment of "M" is "M'=\mathbb{R} \backslash M" , if we show "M'" is open in "\mathbb{R}" we will be done.

Let, "m \in M' \implies m \notin M" ,thus as defined in "(\clubs)" , "B(m,r_m)" and considering "(\spades)" implies



"M'=\cup_{m \in M'} B(m,r_m)"

Hence, "M'" is open in "\mathbb{R}" as arbitrary union of open set is open set.Therefore, "M" is closed in "\mathbb{R}" .

We are done.